### v1

# 1.3in1.3in1in1in

Abstract.
For a cubic fourfold $Y$ with associated Fano variety of lines $F$, we establish several properties of the finite-degree 16 self-rational map $\psi\colon F\dashrightarrow F$ introduced by Voisin. We begin by analyzing the singularities of the nodal quintic with 16 nodes associated with a general line under the specialization to a line in the branch locus of $\psi$. This approach reveals that the ramification of the natural resolution of indeterminacy of $\psi$ is simple. The main part of the paper focuses on the intriguing interplay between $\psi$ and the fixed locus of the antisymplectic involution on the LLSvS variety $Z$, examined via the degree 6 Voisin map $F\times F\dashrightarrow Z$. As an application, we show that the monodromy of $\psi$ is maximal.
Key words and phrases:
Irreducible symplectic varieties, cubic fourfolds
2020 Mathematics Subject Classification:
32J27 (primary), 32S15 (secondary).
1. Introduction
Let $Y$ be a smooth cubic fourfold and $F$ be its (Fano) variety of lines. The rich geometry of these varieties has attracted great attention from the mathematical community for several reasons. Notably, the Fano variety $F$ is one of the earliest examples of a locally complete family of projective hyperkähler manifolds, and the binome $Y-F$ between a Fano variety and a hyperkähler manifold has inspired numerous other constructions.
One distinguishing feature of $F$ is the degree 16 self-rational Voisin map $\psi\colon F\dashrightarrow F$. Remarkably, $\psi$ is the only known finite, non-birational self-map defined on a locally complete family of projective hyperkähler manifolds. Finite maps such as $\psi$ are particularly subtle to study as, unlike birational maps, they cannot be detected through their action on the second cohomology group and lack a comprehensive classification result as Hodge Torelli theorem that would allow precise control over their behavior. Since its construction [Voi04], this map has continued to attract significant interest in the mathematical community, as demonstrated by the numerous works on the subject[Ame09, GK23b, GK23a, GK22].
In this paper, we focus on the monodromy of the Voisin map $\psi$. The monodromy group is a discrete invariant that encodes the symmetries and intrinsic geometry of finite maps. Determining the monodromy group of branched coverings is a classical problem, originating with Jordan in the 1870s. This topic was revitalized by Harris, who provided a modern framework by proving that the monodromy and Galois groups coincide, and developed tools to establish when the monodromy group is maximal. Significant progress was made by Vakil, who introduced innovative techniques to study monodromy in Schubert problems on Grassmannians [Vak06].
Building on these developments, we take a step further by investigating linear spaces on a cubic hypersurface of dimension 4. Our main result shows that the monodromy group of the Voisin map $\psi$ is maximal, meaning that it is the full symmetric group on 16 elements. To achieve this, we leverage another Voisin map involving twisted cubics on the cubic fourfold, which unveils a fascinating connection with the fixed locus of the natural anti-symplectic involution on the LLSvS variety.
We now introduce the necessary notation and recall some useful results in order to state our main theorems. Lines on a cubic fourfold fall into two cases: For a general line $L$ the linear space
$\Lambda_{\ell}:=\bigcap_{p\in L}T_{p}Y$
is 2-dimensional and in this case the line is said to be of type I. For special lines, called of type II, the dimension of $\Lambda_{\ell}$ is 3. For the general line we have $\Lambda_{\ell}\cap Y=2\ell+r$ for a line $r\in F$, and one sets $\psi(\ell):=r$. This map has been studied in detail in [Ame09], its indeterminacy locus consists of the lines of type $\operatorname{II}$, which for a general cubic fourfold form a smooth surface $S_{\operatorname{II}}$. Blowing up $S_{\operatorname{II}}$ provides a resolution of the indeterminacy
[…]
Gounelas and Kouvidakis recently computed that the restricion of ${\widehat{\psi}}$ to the exceptional divisor $E$ of the blow-up morphism, which coincides with the ramification divisor of ${\widehat{\psi}}$, is birational onto the image and they posed the question of whether the ramification of ${\widehat{\psi}}$ is simple [GK23a, Theorem B, Remark 3.13].
In Section 2 we study the ramification of ${\widehat{\psi}}$ by consider the projection of $Y$ from a general line $R$ in $Y$. Its resolution is a conic bundle over $\mathbb{P}^{3}$ with discriminant locus a quintic surface $S_{r}$ with exactly 16 nodes. These nodes correspond to the preimage of $r$ under ${\widehat{\psi}}$. Nodal quintic surfaces have been classically studied by Beauville and Catanese [Bea80, Cat81] and more recently studied in [Huy24, BFK+24, CicwYCC+24]. Our first attention is to study the singularity of $S_{r}$ for a special line $r$, which lies in the branch locus of ${\widehat{\psi}}$.
Theorem 1.1 (see Theorem 2.6).
Let $r\in F$ be a general line in the branch divisor of ${\widehat{\psi}}$. Then the quintic surface $S_{r}$ has one singularity of type $A_{3}$ corresponding to the unique line $L$ of type $\operatorname{II}$ over $r$.
As a result of our analysis we affirmatively answer Gounelas and Kouvidakis’ question:
Corollary 1.2 (see Corollary 2.7).
The ramification at the general point of the ramification locus of ${\widehat{\psi}}$ is simple.
In Section 3 we shift our focus to the geometry of twisted cubics. With any smooth cubic fourfold $Y$ not containing a plane Lehn, Lehn, Sorger, and van Straten associated an eight-dimensional hyperkähler variety $Z$ parametrizing families of twisted cubics and their flat degenerations. The variety $Z$ is equipped with a natural antisymplectic involution $\tau$ [Leh15]. Its fixed locus is a smooth Lagrangian submanifold with 2 connected components: one is isomorphic to the cubic fourfold $Y$, the other one, $W$, is of general type [FMOS23] and remains somewhat mysterious. Using a the degree 6 rational map $\varphi\colon F\times F\dashrightarrow Z$, constructed by Voisin [Voi16], we offer an alternative description of $W$. We define the variety $P$ as the closure in $F\times F$ of
$\{(\ell_{1},\ell_{2})\in F\times F:\ell_{i}\text{ are of type I,\quad$\ell_{1}\not=\ell_{2}$,\quad and }\quad\psi(\ell_{1})=\psi(\ell_{2})\},$
which is birational to an irreducible component of the self-product of ${\widehat{F}}$ over $F$.
Theorem 1.3 (see Theorem 3.7).
The variety $P$ is mapped onto $W$ under the Voisin map $\varphi\colon F\times F\dashrightarrow Z$.
In Section 4 after revising the basic notions of monodromy we tackle the study of ${\widehat{\psi}}$. Despite extensive study and numerous works concerning ${\widehat{\psi}}$, for example about its entropy [Ame09], many of its properties remain elusive. Through an investigation of the restriction of the map $\varphi$ to the variety $P$ we prove the following
Theorem 1.4 (see Theorem 4.4).
The monodromy group of $\psi$ is the entire symmetric group on 16 elements.
Acknowledgments
This project began long time ago, and over the years we benefited from conversations with many people. It is our pleasure to thank everybody who expressed interest and shared their point of view, especially Enrico Fatighenti, Frank Gounelas, Christian Lehn, Emanuele Macrì, Giovanni Mongardi, Alan Thompson and Yilong Zhang.
Franco Giovenzana was funded by Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) Projektnummer 509501007, and partially supported by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (ERC-2020- SyG-854361- HyperK). All authors are members of INdAM GNSAGA.
2. The ramification of the Voisin map is simple
In this section we recall basic facts abour lines on cubic fourfolds and various properties of the Voisin map on $F$. Then we move on to prove that the ramification of the Voisin map is simple.
Recall that the Gauss map associates to any point of the smooth cubic fourfold $Y\subset{\mathbb{P}}(V)$ its projective tangent space:
$\mathscr{G}\colon Y\to{\mathbb{P}}(V^{\vee}),\ P\mapsto T_{P}Y\simeq{\mathbb{P}}^{4}.$
Clemens and Griffiths distinguished lines on cubic hypersurfaces into two types, we recall here the definition for cubic fourfolds [CG72].
Proposition 2.1.
Given a line $L$ on a smooth cubic fourfold $Y$, either the following equivalent conditions hold:
  1. (1)
    $N_{L|Y}\simeq\mathcal{O}_{L}^{\oplus 2}\oplus\mathcal{O}_{L}(1)$,
  2. (2)
    $\mathscr{G}|_{L}\colon L\to\mathscr{G}(L)$ is 1:1,
  3. (3)
    $\mathscr{G}(L)$ is a smooth conic in ${\mathbb{P}}(V^{\vee})$,
  4. (4)
    $\bigcap_{P\in L}T_{P}Y$ is isomorphic to ${\mathbb{P}}^{2}$;
or the following equivalent conditions hold:
  1. (5)
    $N_{L|Y}\simeq\mathcal{O}_{L}(-1)\oplus\mathcal{O}_{L}(1)^{\oplus 2}$,
  2. (6)
    $\mathscr{G}|_{L}\colon L\to\mathscr{G}(L)$ is 2:1,
  3. (7)
    $\mathscr{G}(L)$ is a line in ${\mathbb{P}}(V^{\vee})$,
  4. (8)
    $\bigcap_{P\in L}T_{P}Y$ is isomorphic to ${\mathbb{P}}^{3}$.
Definition 2.2 ([CG72, Definition 6.6]).
Given a line $L$, we say that $L$ is a line of type I if the equivalent conditions $(1)-(4)$ hold, whereas if $(5)-(8)$ hold we say that $L$ is a line of type II. We set $\Lambda_{L}:=\cap_{P\in L}T_{P}Y$.
We record here these elementary observations for future reference.
Remark 2.3.
In case $\ell\in F$ is of type $\operatorname{II}$, then the line $\mathscr{G}(L)\subset{\mathbb{P}}(V^{\vee})$ is the projective dual of $\Lambda_{L}\subset{\mathbb{P}}(V)$. In case of a line of type I, $\mathscr{G}(L)$ spans a ${\mathbb{P}}^{2}$ in ${\mathbb{P}}(V^{\vee})$ which is dual to $\Lambda_{L}$.
Remark 2.4.
[Huy23, Remark 2.2.2] Let $Y=V(F)\subset{\mathbb{P}}^{5}$ be a smooth cubic fourfold, let $L$ be a line in $Y$, then $L$ is of type II if and only if the partial derivatives $\partial_{0}F|_{L},...,\partial_{6}F|_{L}$ span a vector space of dimension 2 in $H^{0}(L,\mathcal{O}_{\mathbb{P}^{5}}(2))$.
Voisin observed that if $L$ is a line of type I, then $\Lambda_{L}\cap Y$ is a cubic curve, which, as it contains $L$ with multiplicity 2, consists of $L$ and a residual line $R$. As the general line is of type I, one defines the rational map
$\psi\colon F\dashrightarrow F,\quad\ell\mapsto r.$
This map has been studied in [Voi04] and [Ame09], where it is proven to be finite of degree 16. Lines of type II form a surface $S_{\operatorname{II}}$ in $F$, and blowing up $F$ in this surface resolves the indeterminacy of $\psi$:
[…]
The blow up ${\widehat{F}}$ can be identified as the closure of the graph [GK22, Lemma 4.1],[Huy23, Remark 2.2.19].
$F\dashrightarrow\operatorname{Gr}(3,6),\ \ell\mapsto\Lambda_{\ell}.$
Elements in the exceptional locus are then just pairs $(\ell,\Xi)$, where $L$ is a line of type II and $\Xi$ is a projective plane such that $L\subset\Xi\subset\Lambda_{\ell}$. For such 2-dimensional spaces $\Xi$ we have that $\Xi\cap Y=2\ell+r$ and we shall say that $r$ is residual to $\ell$.
Given a general line $r\in F$, we consider the diagram
[…]
where $\pi_{R}$ is the projection from $R$ and ${\widetilde{Y}}$ is the blow-up of $Y$ in $R$. The morphism ${\widetilde{\pi}}$ is a conic bundle with discriminant a quintic surface $S_{r}$, whose singular locus consists of 16 nodes. The 16 nodes correspond to the preimage of $R$ under ${\widehat{\psi}}$. Indeed, if $p_{i}$ denote the nodes of $S_{R}$ for $i=1,..,16$, then $L_{i}:=p({\widetilde{\pi}}^{-1}(p_{i}))$ are the 16 lines for which $\psi(L_{i})=R$ (see [Huy23, § 6.4.5] for an account on the various results in the literature about this surface).
Let $E$ be the exceptional divisor of the blow-up morphism ${\widehat{F}}\to F$. As $F$ has trivial canonical bundle, the divisor $E$ coincide with the ramification locus of the map ${\widehat{\psi}}$.
Theorem 2.5 (Gounelas-Kouvidakis,[GK22]).
The restriction ${\widehat{\psi}}|_{E}\colon E\to F$ is generically 1-to-1 onto the image.
In other words, over the general point $r$ in the branch divisor of ${\widehat{\psi}}$, there is exactly one point of ramification, i.e. there exists exactly one line $L$ of type II with residual $R$, meaning that ${\widehat{\psi}}(\ell,\langle L,R\rangle)=r$.
Theorem 2.6.
Let $r\in F$ be a general line in the branch divisor of ${\widehat{\psi}}$. Then the quintic surface $S_{r}$ has exactly 14 singularities of type $A_{1}$ and exactly 1 singularity of type $A_{3}$ corresponding to the unique line $L$ of type $\operatorname{II}$ with residual line $R$.
Proof.
Let $r$ be a general line in the branch locus of ${\widehat{\psi}}$ and let $\ell$ be the unique line of type $\operatorname{II}$ with residual $r$. We may assume that $R$ is of type I and that is given as the zero set $V(x_{2},x_{3},x_{4},x_{5})$. Following Clemens and Griffiths [CG72, equation 6.9] we write the equation of the cubic fourfold as
(1) $x_{4}x_{0}^{2}+x_{5}x_{0}x_{1}+x_{3}x_{1}^{2}+x_{0}Q_{0}+x_{1}Q_{1}+P=0,$
where $Q_{0},Q_{1}$ and $P$ are homogeneous polynomials in the variables $x_{2},x_{3},x_{4},x_{5}$ of degree 2, 2, and 3 respectively. The quintic surface $S_{r}\subset V(x_{0},x_{1})\simeq\mathbb{P}^{3}$ is given by the determinant of the matrix
$M:=\begin{pmatrix}P&Q_{0}&Q_{1}\\ Q_{0}&x_{4}&x_{5}\\ Q_{1}&x_{5}&x_{3}\end{pmatrix}.$
We assume that the line $L$ is given by $V(x_{1},x_{2},x_{4},x_{5})$ [^1] Notice that with this assumption the line $L$ is not contained in the plane tangent to $R$ is $\Lambda_{r}=V(x_{3},x_{4},x_{5})$, would have been in contradiction with $R$ being residual to $L$., then $p:=(1:0:0:0:0:0)$ is the intersection of $L$ and $R$, and $q:=(0:0:0:1:0:0)$ is the point of intersection with $\mathbb{P}^{3}$.
We translate these assumptions in conditions on the coefficients of $Q_{0},\ Q_{1}$, and $P$. For this, we introduce the notation $Q_{0}=\sum a_{I}\underline{x}^{I}$, $Q_{1}=\sum b_{I}\underline{x}^{I}$ and $P=\sum c_{I}\underline{x}^{I}$.
The cubic fourfold $Y$ contains the line $L$ if $F(x_{0},0,0,x_{3},0,0,0)$ is the zero form in $\mathbb{C}[x_{0},x_{3}]$, hence we get the conditon
(2) $a_{33}=c_{333}=0.$
The fact that the line $R$ is residual to $L$, i.e. $\langle L,R\rangle\cap Y=2L+R$, translates into the further condition
(3) $b_{33}=0.$
By Remark 2.4 the line $L$ is of type II can be rephrased in terms of the derivatives $\partial_{i}F|_{L}$, and we get the additional condition
(4) $a_{23}c_{335}-a_{35}c_{233}=0.$
To conclude the proof we check with the aid of Macaulay2 that the surface defined by det$(M)$ for polynomials $Q_{0},\ Q_{1},$ and $P$ with coefficients satisfying conditions (2), (3), and (4) has a singularity of type $A_{3}$ at the point $q$. See ancillary file. ∎
Corollary 2.7.
The ramification at the general point of the ramification locus of ${\widehat{\psi}}$ is simple.
This answers the natural question raised in [GK23a, Remark 3.13].
Proof.
Let $b$ be a general point in the branch divisor of $\psi$. Then by Theorem 2.5 the preimage ${\widehat{\psi}}^{-1}(b)$ consists of a finite number of points $p_{1},\ldots p_{k}$ where only $p_{1}$ is of ramification, while ${\widehat{\psi}}$ is étale at the points $p_{2},\ldots p_{k}$. All the points $p_{i}$ correspond to singularities of the quintic surface $S_{b}$: in particular $p_{2},\ldots p_{k}$ are singularities of type $A_{1}$, in contrast $p_{1}$ is a singular point of type $A_{3}$ by Theorem 2.6. The surface $S_{b}$ is specialization of the surface $S_{r}$ for the general line $r\in F$, which has exactly 16 nodes. Thus we deduce that 2 points among these 16 specialize to the singularity of type $A_{3}$, whereas the remaining 14 specialize to the $p_{2},..,p_{k}$, deducing that $S_{b}$ has 14 singularities of type $A_{1}$ apart from the $A_{3}$ singularity. ∎
Let $r$ be any line on the cubic fourfold, we denote by $F_{r}$ the locus of lines meeting $r$
$F_{r}:=\{\ell\in F:L\cap R\not=\emptyset\}.$
If $r$ is general, then $F_{r}$ is a smooth surface [Voi86, §3  Lemma 1]. As an application of our analysis we determine the singularities of $F_{r}$ for a general point $r$ in the branch locus of ${\widehat{\psi}}$.
Corollary 2.8.
Let $r\in F$ be a general line in the branch locus of ${\widehat{\psi}}$, then the surface $F_{r}$ has one singularity of type $A_{1}$ at the point $\ell$ corresponding to the unique line $L$ of type $\operatorname{II}$ with residual line $R$.
Proof.
Let $\pi_{R}\colon Y\dashrightarrow{\mathbb{P}}^{3}$ be the projection from $R$. Then there is a map
$f\colon F_{r}\dashrightarrow S_{r},\quad\ell\mapsto\pi_{R}(L).$
The map clearly extends on the point $r$ by mapping $r$ to $\pi_{r}(\Lambda_{r})$, when the line $r$ is of type $I$. The map $f$ is 2-to-1 and can be regarded as the quotient of an involution, which, in case $r$ is general, it has 16 fixed points mapped to the 16 singular points of $S_{r}$. If $r$ is as in the hypothesis, then $S_{r}$ is has 14 singularities of type $A_{1}$ and one singularity of type $A_{3}$ corresponding to $\ell$. All the nodes get resolved in the 2-to-1 cover $F_{r}$, while the $A_{3}$ singularity gives an $A_{1}$ singularity, hence the claim. ∎
3. On a lagrangian submanifold of $Z$ and singular cubic surfaces
Given a smooth cubic fourfold $Y$, we denote by $Z=Z(Y)$ the associated LLSvS variety. It carries the action of an antisymplectic involution which has been studied in [FMOS23] whose fixed locus consists of two connected components, one isomorphic to the cubic fourfold, which we hence denote $Y$, and a second one which we denote $W$. Goal of this section is to study the less understood variety $W$. In order to do that we introduce a subvariety $P$ of $F\times F$ which is related to $W$ via the Voisin map $\varphi:F\times F\dashrightarrow Z$. A key tool in our strategy is to reduce statements to the study of cubic surfaces $S$ which are given by intersecting the cubic $Y$ with a projective space $\mathbb{P}^{3}$ and consider lines and twisted cubics on $S$.
3.1. The LLSvS variety $Z$
Given a smooth cubic fourfold $Y$ not containing any plane, Lehn, Lehn, Sorger and van Straten constructed in [LLSvS17] a hyperkähler variety $Z$ of dimension 8, which parametrizes families of twisted cubics on $Y$ and their flat degenerations.
As any twisted cubic has $13t+$ 1 as Hilbert polynomial, the starting point is to consider the compactification $\operatorname{Hilb}^{gtc}(Y)$ of the space of twisted cubics inside the Hilbert scheme $\operatorname{Hilb}^{3t+1}(Y)$. Curves parametrized by this variety are called generalised twisted cubics. The variety $\operatorname{Hilb}^{gtc}(Y)$ is smooth, of dimension 10 and admits a morphism to the Grassmannian $G:=\operatorname{Gr}(4,6)$, which maps any generalised twisted cubic to its linear span. This morphism factors through a ${\mathbb{P}}^{2}$-fibration onto a variety $Z^{\prime}$ which has a generically finite morphism to the Grassmannian. The hyperkähler projective manifold $Z$ is obtained by $Z^{\prime}$ by contracting the divisor parametrising linear systems of curves which are not arithmetically Cohen-Macaulay. It comes with the natural polarisation $\mathcal{L}$ induced by the map to the Grassmannian $G$. It has dimension 8 and the image of the divisor under the contraction is a Lagrangian subvariety isomorphic to the cubic fourfold $Y$.
Remark 3.2.
The construction above of the variety $Z^{\prime}$ is performed relatively over the Grassmannian $G$, so that can be carried out for any cubic surface $S$, thus we shall and will use the variety $Z^{\prime}(S)$ for any cubic surface $S$. It clearly comes equipped with the ${\mathbb{P}}^{2}$-fibration $\mathrm{Hilb}^{gtc}(S)\to Z^{\prime}(S)$.
The hyperkähler manifold $Z$ has a biregular antisymplectic involution $\tau\colon Z\to Z$ as pointed out in [Leh15]. It can be defined as follows: Let $C\subset Y$ be a generalised twisted cubic and let $Q$ be any quadric (surface) in the linear span $\langle C\rangle$, which contains $C$. The intersection $Q\cap Y$ is a sextic which contains $C$, the residual cubic is another generalised twisted cubic $C^{\prime}$. Antisymplectic involutions of K3[n]-type manifolds and this particular example have been studied in [FMOS22, FMOS23].
Theorem 3.3 ([FMOS23, Theorem 1.4]).
The fixed locus of the antisymplectic involution has two connected components, one of which is the lagrangian subvariety $Y$. The other component $W$ is of general type, in particular $\omega_{W}=3\mathcal{L}|_{W}$, where $\mathcal{L}$ is the primitive polarization on $Z$.
Voisin constructed a degree 6 rational map $\varphi\colon F\times F\dashrightarrow Z$ in this way [Voi16]. Given a general pair of skew lines $(\ell_{1},\ell_{2})$ and a point $p\in L_{1}$, then the union of $\ell_{1}$ and the quadric, which is residual to $\ell_{1}$ in the intersection $\langle L_{1},\ p\rangle$, is a generalized twisted cubic. The indeterminacy locus, the resolution of it and the branch divisor of its resolution have been studied in [Mur20, Che21, Gio20].
The involution $\tau$ fits in the following commutative diagram:
${F\times F\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\sigma}$$\scriptstyle{\varphi}$${Z\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\tau}$${F\times F\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\varphi}$${Z,}$
where $\sigma\colon(\ell_{1},\ell_{2})\mapsto(\ell_{2},\ell_{1})$ switches the factors.
3.4. A variety dominating $W$
We denote by $P$ the closure in $F\times F$ of
$\{(\ell_{1},\ell_{2})\in F\times F:\ell_{i}\text{ are of type I, $\ell_{1}\not=\ell_{2}$ and }\psi(\ell_{1})=\psi(\ell_{2})\}.$
It turns out to have a central rôle in our study: The following intriguing lemma is the starting point which allows us to relate properties of $\psi$ and $\phi$.
Lemma 3.5.
Let $(\ell_{1},\ell_{2})\in P$ be a general pair of lines, satisfying the conditions
  1. (1)
    $\ell_{1}$ and $\ell_{2}$ are skew,
  2. (2)
    $\ell_{1}$ and $\ell_{2}$ are both of type I,
  3. (3)
    and $\psi(\ell_{1})=\psi(\ell_{2})$.
Then the linear span of $L_{1}$ and $L_{2}$ intersects $Y$ in a Cayley cubic surface.
Proof.
We claim that the linear span $\Lambda:=\langle L_{1},L_{2}\rangle$ intersects $Y$ in a cubic surface with two singular points on $L_{1}$. We denote by $r$ the residual line $\psi(\ell_{1})$, then $R$ and $L_{1}$ are coplanar and we consider the line $T$ in ${\mathbb{P}}(V^{*})$ of hyperplanes which contain $\Lambda$.
As both $T$ and ${\mathscr{G}}(L_{1})$ lie in the 2 dimensional linear space $\langle L_{1},R\rangle^{*}\subset{\mathbb{P}}(V^{*})$, the line $T$ intersects the smooth conic ${\mathscr{G}}(L_{1})$ in two points, which correspond to the image under the Gauss map of two singular points of $S:=\Lambda\cap Y$ lying on the line $L_{1}$.
Repeating the argument for $L_{2}$ we find two singular points of $S$ on $L_{2}$ and we deduce that the cubic surface $S$ is Cayley by [Dol12, Corollary 9.2.3]. ∎
Following the construction of the Voisin map, one gets the following
Lemma 3.6.
Let $L_{1}$ and $L_{2}$ be two lines that satisfy the general conditions:
  1. (1)
    $L_{1}$ and $L_{2}$ are of type I,
  2. (2)
    $L_{1}$ and $L_{2}$ are skew.
If $\psi(\ell_{1})=\psi(\ell_{2})=:r$ then $\varphi(\ell_{1},\ell_{2})=\varphi(\ell_{2},\ell_{1})$.
We report the following theorem which is a key result for our study, but we defer its proof to later, because it relies on some observations conerning Cayley cubics.
Theorem 3.7.
The Voisin map $\varphi\colon F\times F\dashrightarrow Z$ restricts to a dominant rational map $\varphi|_{P}\colon P\dashrightarrow W$.
The regular map $W\to G$ is birational onto the image, which consists of the projective spaces that intersect $Y$ in a Cayley cubic or a degeneration of it.
In the following we analyze the maps $P\dashrightarrow W\to G$ in greater detail. In order to do so, we describe singular cubic surfaces with focussing on lines on them.
3.8. Cayley cubics and their degenerations
The stratification of the space of cubic surfaces given by the type of singularities has been studied by Bruce in [Bru80]. Here we are interested only in cubic surfaces with exactly 4 singular points of type $A_{1}$, called Cayley cubics, and their degenerations. From his analysis degenerations of Cayley cubics can be summed up by the following diagram:
[…]
where ${\widetilde{E}}_{6}$ denotes the cubics with one simple elliptic singularity and $X_{i}$ are non-normal cubic surfaces.
In the following we focus on lines and twisted cubic on singular cubic surfaces to study the maps $\varphi$ and $\psi$. We notice that the map $\psi$ is in general not defined on $Z(S)$ for a singular cubic surface $S$ because a line $L$ need not have a plane tangent at every point. However, if $\pi\subset\langle S\rangle$ is tangent at every point of $L\subset S$ with $\pi\cap S=2L+R$ for a line $R$, we say that the line $R$ is residual, and that $L$ is triple if $R=L$. Generalised twisted cubics on singular cubic surfaces have already been carefully analysed in [LLSvS17, § 2], which we closely follow. Every cubic surface $S$ with at most double rational points has a minimal resolution ${\widetilde{S}}\to S$ with ${\widetilde{S}}$ isomorphic to the blow-up of $\mathbb{P}^{2}$ in six points. The orthogonal $\Gamma:=K_{{\widetilde{S}}}^{\perp}\subset H^{2}({\widetilde{S}},\mathbb{Z})$ is isomorphic to the negative definite root lattice $E_{6}$. We denote by $R_{0}$ the roots in $\Gamma$ which are effective, they correspond to irreducible components of exceptional curves for the resolution. We also denote by $W(R_{0})$ the subgroup of the Weyl group of $E_{6}$ generated by reflections in the effective roots $R_{0}$. The moduli space of generalised twisted cubics on $S$ with reduced structure is then given by a union of projective planes [LLSvS17, Theorem 2.1]:
$\mathrm{Hilb}^{gtc}(S)_{red}=(R/W(R_{0}))\times{\mathbb{P}}^{2}.$
In particular, linear systems of generalised twisted cubics on $S$ are in bijection with the orbits of $R$ under the action of $W(R_{0})$. These correspond to the set $Z^{\prime}(S):=a^{-1}(E)$ where $E\in G$ is the linear span $\langle S\rangle$. Along this bijection families of arithmetically Cohen-Macaulay (aCM) curves correspond to orbits containing an effective root. By the definition of $\tau$ it is clear that it restricts to an action on twisted cubics contained on a cuibic surface, and indeed admits a nice description in this setting: As observed in [Leh15] it is indeed given by the association $\alpha\mapsto-\alpha$ for the roots $\alpha\in R$.
We defer to Appendix 5 details about these singular surfaces, and recall here only the main facts we need to analyse the the Voisin maps.
We also introduce a piece of notation. We denote by $G_{4A1}$ the closure in the Grassmanian $G$ of
$\{E\in G:E\cap Y\mbox{ is a singular cubic surface of type }4A_{1}\}.$
Analogously, we define the loci $G_{2A_{1}A_{3}}$, $G_{A_{1}A_{5}}$, $G_{{\widetilde{E}}_{6}}$, and for the non-normal surfaces $G_{X_{6}}$, $G_{X_{7}}$, $G_{X_{8}}$, $G_{X_{9}}$.
Singular cubic surfaces of type $4A_{1}$
Any Cayley cubic, i.e. a surface with exactly 4 $A_{1}$ singular points, contains exactly 9 lines whose schematic representation resembles a tetrahedron with vertices at the singular points. We name edges the 6 of them connecting 2 singular points. In particular, if $P_{1},P_{2},P_{3},P_{4}$ are the singular points, we denote $E_{ij}$ the line containing $P_{i}$ and $P_{j}$. Each of the other 3 lines intersect exactly 2 edges, we denote $J_{ij,kl}$ the line joining the edges $E_{ij}$ and $E_{kl}$ for $\{i,j,k,l\}=\{1,2,3,4\}$. See Figure 1(a).
[…]
(a) Lines on a singular cubic surface of type $4A_{1}$
[…]
(b) Lines on a $2A_{1}A_{3}$ surface
Figure 1.
We denote by $K$ the linear system of generalised twisted cubics containing the reducible elements given by the lines through a singular point
$K=[E_{ij}\cup E_{ik}\cup E_{il}]=[E_{ji}\cup E_{jk}\cup E_{jl}]\in Z^{\prime}_{aCM}(S).$
Lemma 3.9.
Opposite edges map to $K$ under $\varphi$, that is:
$\varphi(E_{ij},E_{kl})=K\quad\text{ for indices such that }\{i,j,k,l\}=\{1,2,3,4\}.$
Lemma 3.10.
Let $S$ be a Cayley cubic. Then
#$\#Z^{\prime}_{aCM}(S)=13\quad\mbox{ and }\quad\#Z^{\prime}(S)^{\tau}=1.$
Proof.
The first equality follows from [LLSvS17, Table 1]. The second one can be deduced by the description of $\tau$ in terms of roots. ∎
Proof.
(of Theorem 3.7) By Lemma 3.5 the general point $(\ell_{1},\ell_{2})$ in $P$ corresponds to opposite edges on a Cayley cubic surface. By Lemma 3.9 its image $\varphi((\ell_{1},\ell_{2}))$ is an aCM twisted cubic fixed by the involution on $Z$, i.e. a point in $W$. In particular, $\varphi$ restricts to a dominant map $\varphi:P\dashrightarrow W$. Finally, the morphism $W\to G$ is 1:1 because by Lemma 3.10 given a projective space $\Lambda\simeq\mathbb{P}^{3}$ which intersects $Y$ in a Cayley cubic $S$, $K$ is the unique linear system of $aCM$ generalised twisted cubics in $S$ which is fixed by $\tau$. ∎
Singular cubic surfaces of type $2A_{1}A_{3}$
Let $S$ be a cubic surface with a singular point $P_{2}$ of type $A_{3}$, and 2 singular points $P_{0},P_{3}$ of type $A_{1}$. It contains exactly 5 lines as depicted in Figure 1(b): the line $R_{01}$ through $P_{0},P_{1}$; the line $R_{02}$ through $P_{0},P_{2}$; the line $R_{03}$ through $P_{03}$; the line $L_{1}$ containing the only singular point $P_{2}$, and the line $L_{2}$ contained in the smooth locus.
Proposition 3.11.
Let $S$ be a cubic surface of type $2A_{1}A_{3}$. Then:
$L_{1}\mbox{ is residual to }R_{23},\qquad L_{1}\mbox{ is residual to }R_{02},$
$L_{2}\mbox{ is residual to }R_{03},\qquad L_{2}\mbox{ is residual to }L_{1}.$
Moreover, there is no plane in $\langle S\rangle\simeq\mathbb{P}^{3}$ tangent to $S$ at every point of $L_{2}$. In particular, there are no triple lines on a surface with $2A_{1}A_{3}$ singularities.
Remark 3.12.
If $S$ is a cubic surface of type $2A_{1}A_{3}$ given by intersecting a smooth cubic fourfold $Y$ with a $\mathbb{P}^{3}$, then one can conclude there are no triple lines on S. Indeed, if all the lines are of type I, Proposition 3.11 can be restated as:
$\psi(R_{23})=\psi(R_{02})=L_{1},\text{ and }\psi(R_{03})=\psi(L_{1})=L_{2}.$
If a line, e.g. $R_{23}$, is of type II one can conclude that ${\widehat{\psi}}(R_{23},\pi)=L_{1}$ where $\pi=\langle R_{23},L_{1}\rangle$.
Lemma 3.13.
Let $S$ be a cubic surface of type $2A_{1}A_{3}$. Then
#$\#Z^{\prime}_{aCM}(S)=5\quad\mbox{ and }\quad\#Z_{aCM}^{\prime}(S)^{\tau}=1.$
Proof.
The first equality follows from [LLSvS17, Table 1]. The second one can be deduced by the description of $\tau$ in terms of roots. ∎
Singular cubic surfaces of type $A_{1}A_{5}$
Let $S$ be a cubic surface with an $A_{1}$ and an $A_{5}$ singularity. It contains exactly two lines, say $L_{1}$ passing through both singular points, and $L_{2}$ which intersects the singular locus in the $A_{5}$ point.
Proposition 3.14.
Let $S$ be a cubic surface of type $A_{1}A_{5}$. Then
$L_{2}\text{ is triple, and }L_{2}\text{ is residual to }L_{1}.$
Lemma 3.15.
Let $S$ be a cubic surface of type $A_{1}A_{5}$. Then
#$\#Z^{\prime}_{aCM}(S)=\#Z_{aCM}^{\prime}(S)^{\tau}=1.$
Proof.
Clear from [LLSvS17, Table 1]. ∎
3.16. Simple elliptic cubic surfaces
Any cubic surface $S$ with a simple elliptic singularity is a cone over a planar elliptic curve $E$. Let $v$ be the vertex. The Fano variety $F(S)$ of $S$ can be naturally identified with the embedded elliptic curve $E$, indeed to any point $P$ of $E$ corresponds the line through $P$ and $v$. The scheme of generalised twisted cubics is isomorphic to $\operatorname{Sym}^{3}(E)$ and the summation map to $E$ is a ${\mathbb{P}}^{2}$-bundle giving the fibration $\operatorname{Hilb}^{gtc}(S)\to Z^{\prime}(S)$ [LLSvS17, Proposition 2.7]. A generalised twisted cubic corresponds to the union of 3 lines, which join the vertex with 3 points on $E$. If these are collinear, then they are the support of a nCM curve with an embedded point at the vertex. Via the isomorphism $F(S)\simeq E$, the residual line can be expressed in terms of the group law of $E$, and the analogue of $P$ for lines on the surface $S$
$P(S):=\{(p_{1},p_{2})\in F(S)\times F(S)\mid\mbox{ $p_{1}$ and $p_{2}$ have the same residual}\}$
admits a nice description:
Proposition 3.17.
Let $S$ be a simple elliptic cubic surface. Then
  • the residual to the line $p\in E\simeq F(S)$ is given by $-2p$;
  • $P(S)=\{(p,p+\xi)\mid\xi\in E[2]\}$;
  • the triple lines on $S$ are given by the 3-torsion points $E[3]\subset E$. In particular, every simple elliptic cubic surface has 9 triple lines.
Proposition 3.18.
Let $Y$ be a general cubic fourfold, and let $\ell_{1}$ and $\ell_{2}$ be general lines of type $I$ such that $\psi(\ell_{1})=\psi(\ell_{2})=\ell_{2}$. Then $\langle E_{\ell_{1}},E_{\ell_{2}}\rangle$ is a 3-dimensional space which cuts $Y$ in a cubic surface of type either $A_{1}A_{5}$, or ${\widetilde{E}}_{6}$.
Proof.
By Lemma 3.5 a general point of $P$ consists of a pair of lines whose linear span intersects $Y$ in a Cayley cubic. Thus, if $\ell_{1}$ and $\ell_{2}$ are as in the hypothesis, then the 3 dimensional linear space $\langle E_{\ell_{1}},E_{\ell_{2}}\rangle$ intersects $Y$ in a cubic surface $S$ which is a degeneration of a Cayley cubic and contains $\ell_{2}$ as a triple line on the surface. By the analysis of lines on cubic surfaces, $S$ is either of type $A_{1}A_{5}$, ${\widetilde{E}}_{6}$ or non-normal. As by assumption the lines $\ell_{1}$ and $\ell_{2}$ are not of type $\operatorname{II},$ we can exclude the non-normal cases. ∎
We explore now some connection of our study of singular cubic surfaces with results of Gounelas and Kouvidakis [GK22]. As a consequence, we deduce the existence of a surface of type $A_{1}A_{5}$, which is a fundamental little ingredient for the main theorem.
Let $Y$ be a smooth cubic fourfold. The natural map $\gamma\colon\mathbb{L}\to Y$ from the universal family of lines to the cubic fourfold is a fibration in (2,3)-complete intersection curves. Given a point $y\in Y$, its fiber $\gamma^{-1}(y)$ consists of the lines trough the point $y$. Gounelas and Kouvidakis studied the $g^{1}_{3}$ of these curves coming from the rulings of the corresponding quadric and their relation with the Voisin map.
Corollary 3.19.
Let $Y$ be a general cubic fourfold. Then there exists a $\Lambda\in G$ such that $\Lambda\cap Y$ is a cubic surface of type $A_{1}A_{5}$.
Proof.
Given a pair of lines $\ell_{1}$ and $\ell_{2}$ both of type I, with same residual and such that one is triple, then these are naturally contained in the cubic surface $\langle E_{\ell_{1}},E_{\ell_{2}}\rangle\cap Y$, which is either of type $A_{1}A_{5}$ or ${\widetilde{E}}_{6}$. To discern the 2 cases, we reason as follows. A triple line through the point $y\in Y$ corresponds to a ramification point of the $g^{1}_{3}$ of the curves $C_{y}$ [GK22, Lemma 3.10]. As the general triple ramification point of one of the possibly two $g^{1}_{3}$ of the curve $C_{y}$ occurs when $C_{y}$ is smooth, we deduce that the general $\langle E_{\ell_{1}},E_{\ell_{2}}\rangle$ cuts $Y$ in a cubic surface of type $A_{1}A_{5}$. ∎
Corollary 3.20.
Let $Y$ be a general cubic fourfold.
  1. (1)
    For any $\ell\in S_{\operatorname{II}}$ the intersection $E_{\ell}\cap Y$ is a non-normal singular cubic surface.
  2. (2)
    If $\ell\in V\cap S_{\operatorname{II}}$, then $E_{\ell}\cap Y$ is of type $X_{7}$ or worse.
  3. (3)
    If $\ell\in(V\cap S_{\operatorname{II}})_{\operatorname{sing}}$ is a node, then $E_{\ell}\cap Y$ is of type $X_{8}$ or worse. In particular, there are 3780 projective spaces of dimension 3 intersecting $Y$ in a singular surface of type $X_{8}$.
  4. (4)
    For any $\ell\in S_{\operatorname{II}}$ the surface $E_{\ell}\cap Y$ is not of type $X_{9}$.
Proof.
Given a line $L$ of type $\operatorname{II}$, the cubic surface $E_{\ell}\cap Y$ is singular along the entire $L$. By the study of Gounelas and Kouvidakis a line as in item (2) corresponds to a triple type two line, and one as in item (3) corresponds to a type two line with two triple tangent 2-planes [GK22, comments right above Remark 4.20]. The distinction in cases comes from an elementary analysis of lines on non-normal cubic surfaces (see Appendix 5). The second statement in item (3) follows from [GK22, Theorem B].
For the last item, let $L$ be a line of type $\operatorname{II}$ such that $E_{L}\cap Y$ is a singular cubic surface of type $X_{9}$, that is a cone over a cuspidal cubic curve. Then the lines passing trough the vertex of the cone form a curve, which contains a cuspidal cubic. This is in contradiction with [GK22, Proposition 3.5]. ∎
4. The monodromy group of $\psi$ is maximal
In this section we prove that the monodromy group of ${\widehat{\psi}}$ is maximal. In order to do so, we use that the ramification of ${\widehat{\psi}}$ is simple, as proven in Section 2 and the results of Section 3 about the variety $P$.
Given a generically finite dominant morphism $f\colon X\to Y$ of degree $d$ between irreducible varieties (necessarily of the same dimension), we obtain a degree d extension of function fields $k(X)/k(Y)$, and taking the Galois closure $K/k(Y)$ of this extension, we denote by $\mathrm{Gal}_{f}=\mathrm{Gal}(K/k(Y))$. This agrees with the usual monodromy group $\mathrm{Mon}_{f}$ (see [Har79, SY21]), which is defined as the image in $S_{d}$ of the group of deck transformations of the unramified (i.e., topological) cover $X\setminus f^{-1}(\mathrm{Branch}(f))\to U$, where $U:=Y\setminus\mathrm{Branch}(f)$. Recall the following classical results for $f\colon X\to Y$ and $U$ as above.
Proposition 4.1 ([Har79, p.698]).
If there exists a fibre of $f$ with a unique point of ramification index two and all other points unramified, then $\mathrm{Mon}_{f}\subset S_{d}$ contains a transposition.
Let $X^{[s]}_{U}$ be the complement of the big diagonal in the fibre product of $X_{U}$ $s$-times with itself over $U$. In other words, the fiber of the natural morphism $X^{[s]}_{U}\to U$ consists of $s$ distinct points in the fibre of $f$.
Lemma 4.2 ([SY21, Proposition 2]).
$X^{(s)}_{U}$ is irreducible if and only if $\mathrm{Mon}_{f}$ is an s-transitive subgroup of $S_{d}$.
Thanks to Corollary 2.7 we can readily apply Proposition 4.1 and deduce that the monodromy group $\mathrm{Mon}_{{\widehat{\psi}}}$ contains a transposition. In order to apply Lemma 4.2 we are led to study the irreducibility of ${\widehat{F}}^{(2)}_{U}$ which is strictly related to the variety $P$, we introduced in the previous section. The rest of the section is then devoted to prove the following which is the crucial ingredient to conclude the maximality of monodromy.
Proposition 4.3.
The variety $P$ is irreducible.
We postpone the proof of it and illustrate right away how to get the maximality of monodromy.
Theorem 4.4.
The monodromy of ${\widehat{\psi}}$ is maximal, i.e. $\operatorname{Mon}_{{\widehat{\psi}}}=S_{16}$.
Proof.
From Corollary 2.7 and Proposition 4.1 $\operatorname{Mon}_{{\widehat{\psi}}}$ contains a transposition and thanks to Proposition 4.3 and 4.2 the action of $\operatorname{Mon}_{{\widehat{\psi}}}$ is $2$-transitive. Thus $\operatorname{Mon}_{{\widehat{\psi}}}$ contains all transpositions: Indeed up to renumbering we can assume the transposition $(1,2)\in\operatorname{Mon}_{{\widehat{\psi}}}$. By 2-transitivity there exists $\sigma\in\operatorname{Mon}_{{\widehat{\psi}}}$ such that $\sigma(1)=a$ and $\sigma(2)=b$ for any $a$ and $b$ and we have
$\sigma\circ(1,2)\circ\sigma^{-1}=(a,b).$
Hence $\operatorname{Mon}_{\widehat{\psi}}$ contains any transposition and coincides with the entire symmetric group of 16 elements. ∎
In order to prove the irreducibility of $P$ we take advantage of the map $\varphi\colon P\dashrightarrow W$ and use its various properties we have showed in Section 3, to compute the ramification of $\varphi$ at some special points. Clearly there is an open subset $U_{et}\subset W$ for which the base change $P_{U_{et}}\to U_{et}$ is étale of degree 6 and proper, but as we are interested in the ramification of $\varphi$ we need to study a slightly bigger open subset, thus we provide the following technical proposition.
Proposition 4.5.
There is an open $U_{\mathrm{flat}}\subset W$ such that
  1. (1)
    The pullback $\varphi\colon P_{U_{\mathrm{flat}}}\to U_{\mathrm{flat}}$ is a flat finite morphism.
  2. (2)
    The open set $U_{\mathrm{flat}}$ contains a point in $w\in W$ such that the 3-dimensional linear space $a(w)$ intersects $Y$ in a cubic surface of type $A_{1}A_{5}$.
Proof.
[…]
${\widehat{P}}\setminus{\widehat{\varphi}}^{-1}\left({\widehat{\varphi}}\left(E_{1}\cup E_{\pi}\right)\right)\to Y\setminus{\widehat{\varphi}}(E_{1}\cup E_{\pi})=:U$
is proper, as base change of the proper map ${\widehat{\varphi}}$, and with finite fibers, as we removed the exceptional divisor $E_{1}$). Thus it is a finite morphism. Having removed $E_{\pi}$ the source space is actually a subset of $P$. Hence, we have proven that the restriction of $\varphi$ to $P_{U}\to U$ is a finite morphism.
Moreover, the further restriction of $\varphi$ to a regular locus of $P_{U}$ is a finite morphism of non singular varieties, hence flat by [Liu02, Chapter 4, Remark 3.11]. This open set satisfies then point (1) of the claim.
We now construct explicitly such an open set, for which we show point (2) of the claim. In order to do so, we consider the map $\varphi$ and the auxiliary map $\varphi_{G}$
${P}$${W}$${G.}$$\scriptstyle{\varphi}$$\scriptstyle{\varphi_{G}}$$\scriptstyle{a}$
Their graph closures are going to serve as an explicit resolution of indeterminacy.
${\widehat{P}}:=\mathrm{closure}\{((\ell_{1},\ell_{2}),w)\in P\times W\mid\varphi(\ell_{1},\ell_{2})=w\};$
${\widehat{P}}_{G}:=\mathrm{closure}\{((\ell_{1},\ell_{2}),g)\in P\times G\mid\varphi_{G}(\ell_{1},\ell_{2})=g\}.$
As $\varphi_{G}=a\circ\varphi$, by continuity we have a natural map ${\widehat{P}}\to{\widehat{P}}_{G}$ given by
(5) ${\widehat{P}}\longrightarrow{\widehat{P}}_{G},\quad\left(\ell_{1},\ell_{2},w\right)\longmapsto\left(\ell_{1},\ell_{2},a(w)\right).$
As before, we factor the resolution of indeterminacy ${\widehat{\varphi}}\colon{\widehat{P}}\to W$ in ${\widehat{\varphi}}=\varphi_{2}\circ\varphi_{1}$ using Stein and we denote by $E_{1}$ the exceptional divisor of $\varphi_{1}$ and $E_{\pi}$ the exceptional divisor of $\pi\colon{\widehat{P}}\to P$.
By Corollary 3.19 there exists a point $w\in W$ such that $a(w)\in G$ is a point that intersects $Y$ in a surface $S$ with singularities $A_{1}A_{5}$. By Proposition 3.14 $S$ contains exactly 2 lines $\ell_{1}$ and $\ell_{2}$, that by generality are of type $I$, such that $\psi(\ell_{1})=\ell_{2}=\psi(\ell_{2})$. By the very definition of ${\widehat{P}}_{G}$, its points consist of a pair of lines and a 3-dimensional linear subspace, which contains the 2 lines, hence
${\widehat{\varphi}}_{G}^{-1}\big{(}a(w)\big{)}=\left\{\Big{(}\big{(}\ell_{1},\ell_{2}\big{)},a(w)\Big{)},\ \Big{(}\big{(}\ell_{2},\ell_{1}\big{)},a(w)\Big{)}\right\}.$
Now, as there is just one linear system of arithmetically Cohen-Macaulay generalized twisted cubics on a cubic surface of type $A_{1}A_{5}$ by Proposition 3.15 we get
${\widehat{\varphi}}^{-1}\big{(}w\big{)}=\left\{\Big{(}\big{(}\ell_{1},\ell_{2}\big{)},w\Big{)},\ \Big{(}\big{(}\ell_{2},\ell_{1}\big{)},w\Big{)}\right\}.$
As ${\widehat{\varphi}}^{-1}(w)$ is finite, we have $w\not\in{\widehat{\varphi}}(E_{1})$.
Notice that $\varphi_{G}|_{P}:P\dashrightarrow G$ is defined by the association $(\ell_{1},\ell_{2})\mapsto\langle L_{1},L_{2}\rangle=\langle L_{1},L_{2},R\rangle$ where $r=\psi(\ell_{1})=\psi(\ell_{2})$. Moreover, if both $\ell_{1}$ and $\ell_{2}$ are of type I, then we have $\varphi_{G}((\ell_{1},\ell_{2}))=\langle E_{\ell_{1}},E_{\ell_{2}}\rangle$, hence pairs of distinct lines of type I in $P$ are not in the indeterminacy locus of $\varphi$. Since $Z_{aCM}(S)$ is a singleton for a cubic surface $S$ of type $A_{1}A_{5}$ by Lemma 3.15, the map $\varphi$ is defined at the point $(\ell_{1},\ell_{2})$, thus we conclude that $w\not\in{\widehat{\varphi}}(E_{\pi})$.
Now we are left to prove flatness of $\varphi$ at the points $(\ell_{1},\ell_{2})$ and $(\ell_{2},\ell_{1})$ in $P$. To do so we observe that $(\ell_{1},\ell_{2})$ is a regular point of $P$ as it follows from the lemma below. ∎
Lemma 4.6.
The singular locus of ${\widehat{F}}\times_{{\widehat{\psi}}}{\widehat{F}}$ is contained in $E\times_{{\widehat{\psi}}}E$.
Proof.
Let $(f_{1},f_{2})\in{\widehat{F}}\times_{{\widehat{\psi}}}{\widehat{F}}$ be a point with ${\widehat{\psi}}(f_{1})={\widehat{\psi}}(f_{2})=r$. Then its tangent space is the fiber product over the diagram
[…]
where $T_{f_{1}}{\widehat{F}}$ and $T_{f_{2}}{\widehat{F}}$ have dimension 4 as ${\widehat{F}}$ is a non-singular fourfold. If one among $f_{1}$ and $f_{2}$ is not in the ramification of ${\widehat{\psi}}$, which coincide with the exceptional divisor $E$, then $d{\widehat{\psi}}$ is an isomorphism at that point and the fiber product is a vector space of dimension 4, hence the claim. ∎
Proof.
(of Proposition 4.3) First we are going to prove that $P$ modulo the switching involution $\sigma$ is irreducible. In order to do so, we start with a point $q\in U_{\mathrm{flat}}\cap W_{A_{1}A_{5}}$ which exists by Proposition 4.5. As exactly 2 lines lie on a surface with singularities $A_{1}A_{5}$ (see Proposition 3.14), the preimage $\varphi^{-1}(q)$ consists of two points, say:
$(\ell_{1},\ell_{2})\quad\text{and}\quad(\ell_{2},\ell_{1}).$
Using the involution $\sigma$ on $F\times F$ switching the factors, which clearly restrict to an involution of $P$, we see that $P$ at $(\ell_{1},\ell_{2})$ is locally isomorphic to $P$ at $\sigma((\ell_{1},\ell_{2}))=(\ell_{2},\ell_{1})$. As $\varphi$ is flat, proper and of degree 6 at the point $q$, we conclude that the schematic preimage $\varphi^{-1}(q)$ is a finite scheme with multiplicity 3 both at $(\ell_{1},\ell_{2})$ and $(\ell_{2},\ell_{1})$. We pick now a point $p\in U_{\mathrm{et}}$ close enough to $q$ and focus on the preimages $\varphi^{-1}(p)=\{p_{1},\cdots,p_{6}\}$: If we let $p$ tend to $q$, then 3 points out of the 6 in $\varphi^{-1}(p)$, say $p_{1},p_{2}$ and $p_{3}$, tend to $(\ell_{1},\ell_{2})$ and the other 3 points, i.e. $p_{4},p_{5}$ and $p_{6}$, tend to $(\ell_{2},\ell_{1})$. Moreover, $\sigma$ maps the first triple to the second one, i.e. $\sigma(\{p_{1},p_{2},p_{3}\})=\{p_{4},p_{5},p_{6}\}$, thus $p_{1},p_{2}$ and $p_{3}$ lie in the same connected component of $P$, in other words the quotient of $P$ by the switching involution is path-connected. By Lemma 4.6 we deduce that it is irreducible too.
Hence, it is now sufficient to show that for some point $p_{1}\in P_{U_{et}}$, the points $p_{1}$ and $\sigma(p_{1})$ lie in the same connected component of $P_{U_{et}}$. To do so, we consider one transposition $\tau$ in $\operatorname{Mon}_{\psi}$, whose existence is granted by Proposition 4.1 and Corollary 2.7. If we consider it as an element of deck transformations, this means that there exists a point $r\in F\setminus\operatorname{Branch}({\widehat{\psi}})$, and a loop $\gamma$ centered at $r$ such that $\gamma$ acts on the fiber of $r$ exchanging two points, say $\ell_{1}\mapsto\ell_{2}$ and $\ell_{2}\mapsto\ell_{1}$, and as identity on the remaining other points of the fiber. We now consider the action under deck transformation of $\gamma$ for the natural map $P\dashrightarrow F$ defined by $\psi$. A moment’s thought tells us that under the action of $\gamma$ the point $p=(\ell_{1},\ell_{2})\in P$, with $\psi(\ell_{1})=\psi(\ell_{2})=r$, is mapped to $(\ell_{2},\ell_{1})$. Thus $P$ is irreducible as we claimed. ∎
5. Appendix
In this appendix we study singular cubic surfaces that are degenerations of a cubic surface with exactly 4 singular points of type $A_{1}$. These surfaces are called Cayley.
The stratification of the space of cubic surfaces given by the type of singularities has been studied by Bruce in [Bru80]. From his analysis Cayley cubics can degenerate to a singular surface of type $2A_{1}A_{3}$, which in turn degenerates to a singular surface $A_{1}A_{5}$, which in turn degenerates to surfaces with simple elliptic singularities and then to non-normal ones. Here is the diagram of degenerations:
$4A_{1}\succeq 2A_{1}A_{3}\succeq A_{1}A_{5}\succeq{\widetilde{E}}_{6}\succeq X_{6}\succeq X_{7}\succeq X_{8}\succeq X_{9}$
where ${\widetilde{E}}_{6}$ denotes the cubics with one simple elliptic singularity and $X_{i}$ are non-normal cubic surfaces. We discuss them in details in the subsequent subsections, focussing on lines and generalised twisted cubics and the restrictions of the Voisin maps $\varphi$, and $\psi$ to a cubic surface.
Singular cubic surfaces of type $4A_{1}$
Under a suitable change of coordinate every cubic surface with 4 singular points of type $A_{1}$ can be expressed as zero locus of the following polynomial:
$F=t_{0}t_{1}t_{2}+t_{0}t_{1}t_{3}+t_{0}t_{2}t_{3}+t_{1}t_{2}t_{3}$
Its singular locus consists of the coordinates points
$p_{0}=(1:0:0:0),\quad p_{1}=(0:1:0:0),\quad p_{2}=(0:0:1:0),\quad p_{3}=(0:0:0:1).$
Such a surface can be obtained as the anticanonical model of the blow-up of $\mathbb{P}^{2}$ in special configuration of points. Indeed, let $\widetilde{S}$ be the blow-up of $P^{2}$ in the six points given by the intersection of 4 general lines $L_{0},L_{1},...,L_{3}$. Then the anticanonical linear system of $\widetilde{S}$ maps to a cubic surfaces $S$ and contracts the strict tranforms of $L_{i}$ to 4 singular $A_{1}$ points. The exceptional divisors map to lines in a tetrahedron-like disposition, we refer to them as edges, and denote $E_{ij}$ the line connecting the $i$-th and the $j$-th coordinate point. In particular, the edge $E_{ij}$ is cut by the equations $(t_{k}=t_{l}=0)$ where $\{i,j,k,l\}=\{1,2,3,4\}$. Besides the edges there are 3 more lines on $\bar{S}$ Joining opposite edges; we denote by $J_{ij,kl}$ the line intersecting the edges $E_{ij}$ and $E_{kl}$, where $\{i,j,k,l\}=\{1,2,3,4\}$. These are images of the strict transforms of lines through the blown-up points corresponding to the edges it intersects. For example, the line $J_{01,23}$ is the zero locus $(t_{0}+t_{1}=t_{2}+t_{3}=0)$.
Generalised twisted cubics on singular cubic surfaces have been studied in [LLSvS17, § 2]. Any generalised twisted cubic on an integral normal cubic surface moves in a 2-dimensional linear system and can be described as the image of the pull-back of a line in $\mathbb{P}^{2}$ along the diagram of a determinantal representation of $S$. The determinantal representation for the Cayley cubic described above defines $K\in Z^{\prime}(S)$ which contains as reducible elements of the family, the union of all the edges passing through a fixed vertex of the tetrahedron:
$K=[E_{ij}\cup E_{ik}\cup E_{il}]\in Z^{\prime}(S).$
Lemma 5.1.
Opposite edges map to $K$ under $\varphi$, that is:
$\varphi(E_{ij},E_{kl})=K\quad\text{ for indices such that }\{i,j,k,l\}=\{1,2,3,4\}.$
Proof.
The proof is elementary and can be checked directly by the definition of the Voisin map $\varphi$. For example to compute $\varphi(E_{02},E_{13})$, consider the plane $\pi=(t_{1}=0)$ containing the line $E_{02}$ and the point $p_{\in}E_{13}$. Its intersection with $S$ is given by $\pi\cap S=E_{02}\cup E_{03}\cup E_{23}$, hence $\varphi(E_{02},E_{13})=\left[E_{03}\cup E_{13}\cup E_{23}\right]$ and the claim is shown.
The other cases are completely similar. ∎
Singular cubic surfaces $2A_{1}A_{3}$
Let $S$ be a cubic surface with a singular point $P_{2}$ of type $A_{3}$, and 2 singular points $P_{0},P_{3}$ of type $A_{1}$. It contains exactly 5 lines as depicted in Figure 1(b): the line $R_{01}$ through $P_{0},P_{1}$; the line $R_{02}$ through $P_{0},P_{2}$; the line $R_{03}$ through $P_{03}$; the line $L_{1}$ containing the only singular point $P_{2}$, and the line $L_{2}$ contained in the smooth locus.
Such a surface is given in an opportune coordinate system by the equation
$t_{0}t_{2}t_{3}-t_{1}^{2}t_{3}-t_{0}t_{1}^{2}=0$
Its singular locus consists of the point $P_{2}:=(0:0:1:0)$ of type $A_{3}$ and of the points $P_{0}:=(1:0:0:0)$, $P_{3}:=(0:0:0:1)$ of type $A_{1}$.
The lines are then described by the equations $R_{23}=\{t_{0}=t_{1}=0\},R_{02}=\{t_{1}=t_{3}=0\}$ and $R_{03}=\{t_{1}=t_{2}=0\}$, $L_{1}=(t_{0}=t_{3}=0)$, and $L_{2}=(t_{2}=t_{0}+t_{3}=0)$.
Proposition 5.2.
Let $S$ be a cubic surface of type $2A_{1}A_{3}$. Then:
$L_{1}\mbox{ is residual to }R_{23},\qquad L_{1}\mbox{ is residual to }R_{02},$
$L_{2}\mbox{ is residual to }R_{03},\qquad L_{2}\mbox{ is residual to }L_{1}.$
Moreover, there is no plane in $\langle S\rangle\simeq\mathbb{P}^{3}$ tangent to $S$ at every point of $L_{2}$. In particular, there are no triple lines on a surface with $2A_{1}A_{3}$ singularities.
Proof.
The proof can be checked through straightforward computations. For example, intersecting the surface $S$ with the plane $\pi=(t_{0}=0)$, one gets:
$\pi\cap S$ $=(t_{0}=0)\cap(t_{0}t_{2}t_{3}-t_{1}^{2}t_{3}-t_{0}t_{1}^{2}=0)$
$=(t_{0}=0)\cap(t_{1}^{2}t_{3}=0)=2R_{23}+L_{1},$
which shows that $L_{1}$ is residual to $R_{23}.$
Singular cubic surfaces $A_{1}A_{5}$
A singular cubic surface $S$ with exactly one $A_{1}$ and one $A_{5}$ singularities is given in an opportune of coordinate system by the equation
$t_{3}(t_{0}t_{2}-t_{1}^{2})+t_{0}^{3}=0$
whose singular locus consists of the 2 points $P_{1}:=(0:0:0:1)$ of type $A_{5}$ and $P_{2}:=(0:0:1:0)$ of type $A_{1}$. There are exactly 2 lines on $S$ and they have equations $L_{1}=\{t_{0}=t_{1}=0\}$ and $L_{2}=\{t_{0}=t_{3}=0\}$.
Proposition 5.3.
Let $S$ be a cubic surface of type $A_{1}A_{5}$. Then
$L_{2}\text{ is triple, and }L_{2}\text{ is residual to }L_{1}.$
Proof.
Straightforward computation: The intersection
$\{t_{3}=0\}\cap S=3L_{2}.$
shows the first statement, while
$\{t_{0}=0\}\cap S=L_{1}+2L_{2}$
shows the second one. ∎
The cubic surface $X_{6}$
In a suitable system of coordinates, the surface $X_{6}$ is given by the equation
$t_{0}^{2}t_{2}+t_{1}^{2}t_{3}=0.$
It is a non-normal surface whose singular locus consists of the line $L=(t_{0}=t_{1}=0)$. We record only the following which is necessary for the proof of Corollary 3.20
Proposition 5.4.
The line $L$ is not triple on $X_{6}$.
Proof.
The proof is an elementary computation: Each plane containing $L$ is of the form $\pi=(at_{0}+bt_{1}=0)$ with $(a:b)\in\mathbb{P}^{1}$. By considering the intersection with such a plane $\pi$ one has:
$X_{6}\cap(-t_{0}+bt_{1}=0)=(-t_{0}+bt_{1}=t_{1}^{2}(b^{2}t_{2}+t_{3})=0)=2L+R$
with $R=b^{2}t_{2}+t_{3}=bt_{1}+t_{0}=0$ different from $L$, and
$X_{6}\cap(t_{1}=0)=(t_{1}=t_{0}^{2}t_{2}=0)=2L+R$
with $R=(t_{1}=t_{2}=0)$. This shows that $L$ is not triple. ∎
The cubic surface $X_{7}$
In a suitable system of coordinates, the surface $X_{7}$ is given by the equation
$t_{1}^{3}+t_{2}^{3}+t_{1}t_{2}t_{3}.$
It is a non-normal surface isomorphic to the cone over a nodal planar cubic whose singular locus consists of the line $L=(t_{1}=t_{2}=0)$. We record only the following which is necessary for the proof of Corollary 3.20
Proposition 5.5.
There is a unique plane $\pi$ making $L$ triple, i.e. such that $\pi\cap X_{7}=3L$.
Proof.
The proof is an elementary computation: Each plane containing $L$ is of the form $\pi=(at_{1}+bt_{2}=0)$ with $(a:b)\in\mathbb{P}^{1}$.
By considering the intersection with such a plane $\pi$ one has:
$X_{7}\cap(t_{0}-bt_{1}=0)=(t_{0}-bt_{1}=t_{1}^{2}(bt_{2}+b^{2}t_{3}+t_{1})=0)=2L+R.$
with $R=(t_{0}-bt_{1}=t_{1}+bt_{2}+b^{2}t_{3}=0)$. Also, one has:
$X_{7}\cap(t_{0}=0)=(t_{0}=t_{1}^{3}=0)=3L.$
This shows the claim. ∎
The cubic surface $X_{8}$
In a suitable system of coordinates, the surface $X_{8}$ is given by the equation
$t_{0}t_{1}t_{2}+t_{0}^{2}t_{3}+t_{1}^{3}=0.$
It is a non-normal surface whose singular locus consists of the line $L=(t_{0}=t_{1}=0)$. We record only the following which is necessary for the proof of Corollary 3.20
Proposition 5.6.
There are 2 planes $\pi$ and $\pi^{\prime}$ making $L$ triple, i.e. such that $\pi\cap X_{8}=\pi^{\prime}\cap X_{8}=3L$.
Proof.
The proof is an elementary computation: Each plane containing $L$ is of the form $\pi=(at_{1}+bt_{2}=0)$ with $(a:b)\in\mathbb{P}^{1}$.
By considering the intersection with such a plane $\pi$ one has:
$X_{8}\cap(t_{1}-bt_{2}=0)=(t_{1}-bt_{2}=t_{2}^{2}(b^{3}t_{2}+t_{2}+bt_{3})=0)=2L+R.$
with $R=(t_{0}-bt_{1}=t_{1}+bt_{2}+b^{2}t_{3}=0)$. Also, one has:
$X_{8}\cap(t_{2}=0)=(t_{1}^{3}=t_{2}=0)=3L.$
The planes $\pi=(t_{1}=0)$ and $\pi^{\prime}=(t_{2}=0)$ are hence the only planes as in the claim. ∎